Cyano radicals abstract H atoms from methane at 389.0 K according to the second-order reaction
CN + CH4 → HCN + CH3
Under conditions where [CH4] = 0.493 M is present in great excess (i.e. pseudo-first-order conditions), the concentration of [CN] was observed to decrease from 1.807E-6 M to 2.0167E-12 M in 34.5 ns. Calculate the second-order rate constant for the reaction under these conditions.
Answer #1

Since under excess amount of [CH4 ] the rate law will be

Rate = r = k [ CN- ]1

And it doesn't depends upon CH4 so that

Rate constant for pseudo first order is given by

k = 2.303 / t log { [ A]• / [A] }

Where t is time and [A]• is initial concentration of CN- and [A] is remaining concentration . So,

k = 2.303 / 34.5 x 10^-9 { log { ( 1.807 x 10^-6 / 2.0167 x 10^-12 ) }

k = 0.066 x 10^9 x 5.952 = 3.92 x 10^8 s​​​​​​-1​​​​

in this question ther is a use of first order instead of second order as CH4 is in excess